# Auto-Calibration and Biconvex Compressive Sensing with Applications to Parallel MRI

### v2

@@ Abstract
Keywords: Self-calibration, Compressive sensing, Convex optimization, Random matrices, Parallel MRI.
@@ 1 Introduction
The inverse problem for auto-calibration is generally difficult to solve. In practice, most self-calibration algorithms rely on joint estimation of both the signals of interest and the calibration parameters using standard techniques such as maximum likelihood estimation, alternating minimization, or convex/nonconvex optimization. In this work, we extend the line of work that combines convex optimization and compression sensing in solving various auto-calibration tasks [ling_strohmer_2015, Candès_Eldar_Strohmer_Voroninski_2013, Ahmed_Recht_Romberg_2014, flinth_2017]. A common scheme in such approaches includes first transforming the biconvex constraints to convex constraints using the concept of ‘lifting’ [Candès_Eldar_Strohmer_Voroninski_2013]; then, drawing on ideas from compressive sensing to exploit redundancy/sparsity priors in the signal model and/or the calibration model while being consistent with the convex constraints.
@@
Of particular interest is the case when $\mathbf{x}$ is transform-sparse, a characteristic often observed in real-world signals  [^1] This notion can be easily generalized to approximately transform-sparse.. That is, there is some known sparsifying transformation $\boldsymbol{\Psi}\in\mathbb{C}^{N\times N}$ such that $\boldsymbol{\Psi}\textbf{x}$ is n-sparse with $n<N$. Common choices for $\boldsymbol{\Psi}$ include the well-known DCT basis and the wavelet basis. Despite the sparsity assumption, the available measurements are still fewer than the number of unknowns (i.e., $CL\leq CN+n$). We impose an additional subspace assumption, $\mathbf{s}_{i}=\mathbf{B}\mathbf{h}_{i}$, which can be interpreted as that the calibration parameters all lie in some known subspace spanned by a tall matrix $\mathbf{B}$. Practical choices for such $\mathbf{B}$ can be a low-order polynomial basis or a low-frequency sinusoidal basis. Such choices promote smoothness in the sensing parameters. Altogether, the forward model becomes:
@@
An important observation is that Parallel Imaging (PI) studies can sometimes appear “noisy”. This is largely because the primary objective of PI techniques is to reduce imaging time, which leads to acquiring fewer data points per coil due to the need to reduce imaging time. Consequently, this reduction results in a lower signal-to-noise ratio (SNR). As a result, the SNR of PI sequences is typically lower than that of comparable non-PI sequences [Aja-Fernández_Vegas-Sánchez-Ferrero_Tristán-Vega_2014, Robson_Grant_Madhuranthakam_Lattanzi_Sodickson_McKenzie_2008].
@@
where $R$ denotes the reduction or acceleration factor for undersampling, and $g$ represents a spatially dependent term known as the geometry factor. The term $g$ relates to the number, size, and orientation of the surface coil elements and can be considered a measure of coil separation [Aja-Fernández_Vegas-Sánchez-Ferrero_Tristán-Vega_2014, Robson_Grant_Madhuranthakam_Lattanzi_Sodickson_McKenzie_2008].
@@ 2 Related work
Various auto-calibration problems of similar setups have been explored, including blind deconvolution in [Ahmed_Recht_Romberg_2014], direction-of-arrival estimation as seen in [ling_strohmer_2015], blind deconvolution and demixing [flinth_2017], phase retrieval via matrix completion [Candès_Eldar_Strohmer_Voroninski_2013], among others. The analysis of the auto-calibration problems generally differs in the following two aspects:
@@
Two primary approaches to solving the bilinear optimization problem are as follows: one involves convexifying it through methods such as linearization or lifting [Ahmed_Recht_Romberg_2014, ling_strohmer_2015, Ling_Strohmer_2018, flinth_2017, Candès_Eldar_Strohmer_Voroninski_2013]. The other involves optimization in the natural parameter space using alternating minimization and gradient descent algorithms [Lee_Tian_Romberg_2018, Li_Ling_Strohmer_Wei_2019, Qu_Li_Zhu_2019].
For a general discussion on the issue of injectivity and the principle of identifiability for bi-linear compressive sensing, one may refer to [Kech_Krahmer_2017, Choudhary_Mitra_2013].
@@ 2.1 Related work in auto-calibration and compressive sensing
However, the above nonlinear least squares is a challenging optimization problem since it is highly nonconvex and most of the available algorithms, such as alternating minimization and gradient descent, may suffer from getting easily trapped in some local minima.
@@
Hence, by defining linear operator $\mathcal{A}:\mathbb{C}^{N_{1}\times N_{2}}\rightarrow\mathbb{C}^{m}$, the system of equations becomes
$[y_{1},\cdots,y_{m}]=\mathcal{A}(\mathbf{X}):=\{\mathbf{b}_{i}^{*}\mathbf{X}\mathbf{a}_{i}\}_{i=1}^{m}.$ (6)
@@
An important auto-calibration problem pertaining to sparse signals stems from the task of image reconstruction using randomly coded masks [Bahmani_Romberg_2015, Tang_Recht_2014]. The forward model for this auto-calibration problem is given by:
@@
The SparseLift framework [ling_strohmer_2015] solves the auto-calibration problem in (7) by combining ‘lifting’ and bi-convex compressive sensing. The authors in [ling_strohmer_2015] first transform the biconvex constraint (7) to a linear constraint with respect to $\textbf{X}=\textbf{h}\textbf{x}^{T}$. Exploiting the sparsity in $\mathbf{x}$ and consequently within X, the recovery solves a convex relaxation by minimizing the $\|\mathbf{X}\|_{1}$ subject to the linear constraints. The authors show the recovery of $\mathbf{X}_{0}=\mathbf{h}_{0}\mathbf{x}_{0}^{T}$ is guaranteed with high probability when the sensing matrix $\mathbf{A}$ is either a Gaussian random matrix or with its rows chosen uniformly at random with replacement from the discrete Fourier transformation (DFT) matrix, when the number of measurements $L$ is of the order of $\mathcal{O}(kn\log^{2}(L))$. Note that the case where the signal is only approximately sparse is not analyzed. In a related work, the author in [flinth_2017] extends the results of SparseLift to the case where one observes a summation of such measurements from multiple sparse signals, $\mathbf{y}=\sum\limits_{i\in[r]}\text{diag}(\mathbf{B}\mathbf{h}_{i})\mathbf{A}\mathbf{x}_{i}+\boldsymbol{\omega}$. This extension improves recovery guarantees by solving a $\ell_{1,2}$ minimization by promoting block sparsity in the lifted signals.
The idea of ‘lifting’ also applies to another class of self-calibration problems, notably the blind deconvolution problem [Ahmed_Recht_Romberg_2014, Levin_Weiss_Durand_Freeman_2011]. In the general form of blind deconvolution, one measures the convolution of a signal with an unknown filter. That is, $\textbf{y}=I_{\Omega}(\textbf{s}*\textbf{x})+\textbf{n}$ where both s and x are unknown. Notice that the forward model for our problem in (2) can be written in the form of a convolution $\mathbf{y}=\mathbf{I}_{\Omega}(\mathbf{F}\mathbf{Bh}*\mathbf{F}\boldsymbol{\Psi^{*}}\mathbf{z})+\textbf{n}$ using the convolution theorem. When the Fourier space is fully sampled (i.e. $\mathbf{I}_{\Omega}=\mathbf{I}$ ), the problem is equivalent to recovery from
@@
where $\mathbf{w}$ is also Gaussian. The work by Ahmed et al. [Ahmed_Recht_Romberg_2014] provides a theoretically guaranteed solution to the fully sampled case in (8). Using lifting, the authors solve a nuclear norm minimization problem by exploiting low-rankness in the lifted signal. If the active coefficients of z are known and $\|\textbf{z}\|_{0}=n$, the main theorem of [Ahmed_Recht_Romberg_2014] guarantees a stable recovery when the number of measurements is, up to log factor, of the order $\mathcal{O}(n+k)$ when $\boldsymbol{\Psi}$ is a random Gaussian matrix and B is incoherent. For the more difficult case when the convolution is only subsampled, the authors in [Lee_Li_Junge_Bresler_2017] provide rigorous analysis when both s and x are transform sparse (i.e., there exists $\boldsymbol{\Phi}$ and $\boldsymbol{\Theta}$ such that $\boldsymbol{\Phi}\textbf{x}$ and $\boldsymbol{\Theta}\textbf{s}$ are sparse). It provides an iterative algorithm for robust recovery guarantees with near optimal sample complexity. However, the recovery can only be guaranteed when both the sparsifying transformations (i.e. $\boldsymbol{\Phi}$ and $\boldsymbol{\Theta}$) are random Gaussians. As seen in many theories, subgaussian random matrices provide optimal measurement matrices for compressive sensing [foucart_rauhut_2015]. However, note that for our application case, $\boldsymbol{\Phi}=\boldsymbol{\Psi}\textbf{F}^{-1}$ and $\boldsymbol{\Theta}=\textbf{B}^{*}\textbf{F}^{-1}$, both impose structures due to physical constraints.
Similar to the parallel imaging (i.e., $C>1$ in (2)), from a practical viewpoint, a relevant scenario focuses on self-calibration from multiple snapshots. For example, blind image restoration from multiple filters [Harikumar_Bresler_1999] and self-calibration model for sensors [Balzano_Nowak_2007, Ling_Strohmer_2018, Gribonval_Chardon_Daudet_2012, Bilen_Puy_Gribonval_Daudet_2014]. The goal here is to recover the unknown gains/phases $\mathbf{S}=\text{diag}(\mathbf{s})$ and a signal matrix $\mathbf{X}=[\mathbf{x}_{1},\cdots,\mathbf{x}_{p}]$ from the measurement matrix $\mathbf{Y}=[\mathbf{y}_{1},\cdots,\mathbf{y}_{p}]$ and
@@
which is a linear equation in the concatenated vector $\begin{bmatrix}\mathbf{d}\\ \mathbf{x}_{l}\end{bmatrix}$. Compared to ‘lifting’, an advantage of ‘linearization’ is that the underlying problem dimension remains unchanged. The same idea of linearization can also be found in [Gribonval_Chardon_Daudet_2012, Bilen_Puy_Gribonval_Daudet_2014, Balzano_Nowak_2007, Wang_Chi_2016]. In [Ling_Strohmer_2018], the authors use ‘linearization’ and solve a least-square minimization problem for (9). Furthermore, theoretical recovery guarantee is provided when $\mathbf{A}$ is Gaussian.
As observant readers already may have noticed, the ‘true dimension’ of the problem of recovering $(\textbf{h}_{1},\cdots,\textbf{h}_{c},\textbf{z})$ is $(kC+n)$. To the best of the authors’ knowledge, currently there are no convex optimization schemes which succeed with such few measurements. In this context, the work by Oymak et al. [Oymak_Jalali_Fazel_Eldar_Hassibi_2015] should be mentioned. The authors demonstrated nonconvex procedures (e.g., minimizing a linear combination of $\ell_{0}$, $\ell_{0,2}$ or rank of the lifted matrix) that succeed with high probability, when the number of measurements is (up to log-factors) slightly larger than the information-theoretical limits. However, the theories apply only to measurement ensembles with specific properties, such as matrices with sub-Gaussian rows, subsampled standard basis (e.g. in matrix completion), or quadratic measurements for phase retrieval. Also, these procedures are not efficiently computable. For other structured measurement ensembles, it is worth noting the works by Qu et al. and Ling et al. [Qu_Li_Zhu_2019, Li_Ling_Strohmer_Wei_2019]. These studies address certain blind deconvolution problems with efficient algorithms and theoretical guarantees when the number of measurements is (up to log-factors) slightly larger than the information-theoretical limits. However, to the best of our knowledge, none of these methods directly apply to our measurement ensemble or require additional assumptions that do not align with our application case.
@@ 2.2.1 Relevant problem formulations and the optimization problems
According to linear algebra and the generalized sampling theorem, the maximum reduction factor is equal to the number of channels if E is of full rank. Because these conditions are rarely guaranteed in practice, the possible reduction factor is much lower. In practice, the reduction factor is typically chosen between 2 and 6 [Deshmane_Gulani_Griswold_Seiberlich_2012, Glockner_Hu_Stanley_Angelos_King_2005] for parallel imaging when the number of coils is large. Thus, at the boundary case, one would expect a threshold for the maximum reduction factor and the minimum number of channels necessary to ensure the invertibility of such a linear problem.
@@
As a convex relaxation to $\ell_{0}$ minimization, the following $\ell_{1}$ minimization is often utilized for practical compressive sensing MRI applications. Since Lustig et al. [lustig_donoho_pauly_2007] first demonstrated the compressive sensing approach for MRI reconstruction, it has become one of the most important tools used for modern MR imaging research. The original $\ell_{1}$ norm minimization problem solves the following convex program,
@@
where the signal x is assumed to be transform sparse. In the MRI literature, it has been demonstrated that the discrete cosine transformation, the wavelet transform, and the finite difference transform are the most commonly utilized choices for sparsifying MR images [shin_larson_ohliger_elad_pauly_vigneron_lustig_2014, Ravishankar_Bresler_2015, Chun_Adcock_Talavage_2014]. The choice of such $\boldsymbol{\Psi}$ is related to the field of research in sparsifying transformations [Elad_Milanfar_Rubinstein_2007, Pratt_Kane_Andrews_1969]. More recently, attention has turned to learning the sparsifying transformations from data [Aharon_Elad_Bruckstein_2006, Ravishankar_Bresler_2012].
In terms of combining CS with parallel imaging, the first applications were demonstrated in [Liang_Liu_Wang_Ying_2009, Bo_Liu_Yi_Ming_Zou_Leslie_Ying_2008, Ji_Chen_Zhao_Tao_Lang_2008, Wu_Millane_Watts_Bones_2010]. They can be seen as a direct extension of (13) to the parallel case,
$\min\limits_{\textbf{x}}\|\boldsymbol{\Psi}\textbf{x}\|_{1}+\alpha\|\textbf{x}\|_{TV}\quad\text{subject to }\textbf{E}\textbf{x}=\textbf{y},$ (14)
@@
Note that both the convex relaxations in (13) (14) and nonconvex problem in (12) assume that the coil sensitivities are known. In practice, coil sensitivities are typically determined either through a separate pre-scan [pruessmann_weiger_scheidegger_boesiger_1999, mathematics_and_physics_of_emerging_biomedical_imaging_1996, Liang_Lauterbur_2000] or by interpolating from a fully sampled region of k-space (Fourier space) [griswold_jakob_heidemann_nittka_jellus_wang_kiefer_haase_2002, McKenzie_Yeh_Ohliger_Price_Sodickson_2002, Morrison_Jacob_Do_2007]. However, both methods have inherent drawbacks. The pre-scan approach is susceptible to motion artifacts, environmental fluctuations, and requires additional acquisition time. The interpolation method, while faster, may introduce errors due to imperfect k-space sampling or interpolation techniques. Moreover, both approaches may necessitate the use of filtering and kernel selection to obtain reasonable estimates of the coil sensitivities, which can lead to variations in the resulting reconstructions.
@@
Prior knowledge about the object and the coil sensitivities must be incorporated to render reasonable solutions. As formulated in Joint Image Reconstruction and Sensitivity Estimation in SENSE (JSENSE) [Ying_Sheng_2007] and Regularized Nonlinear Inversion (NLINV) [Uecker_Hohage_Block_Frahm_2008], the smoothness in the coil sensitivities is exploited. In JSENSE, the coil sensitivities are assumed to be expanded by a low-order polynomial basis (i.e. B) and the following bi-convex optimization problem is solved via alternating minimization,
@@
where $\mathbf{W}\in\mathbb{C}^{N\times N}$ is an invertible preconditioning matrix that penalizes high frequencies, thereby enforcing smoothness in the coil sensitivities. Other methods to enforce smoothness in the coil include expanding the coil sensitivities using low-order polynomial or sinusoidal basis functions, a method first introduced in the original SENSE paper [pruessmann_weiger_scheidegger_boesiger_1999] and later adopted in the analytical phantom package [Guerquin-Kern_Lejeune_Pruessmann_Unser_2012].
An extension of the NLINV (16) framework is the ENLIVE [Holme_Rosenzweig_Ong_Wilke_Lustig_Uecker_2019], which solves the following relaxed joint nonconvex optimization problem over $k$ sets of measurements,
@@
Of particular interest to us is that ENLIVE can be related to a nuclear norm minimization problem that exploits the low-rankness in a ‘lifted’ domain. Specifically, by defining the matrix $\mathbf{y}=\begin{bmatrix}\mathbf{y}_{1}&\cdots&\mathbf{y}_{C}\end{bmatrix}$, $\mathbf{U}=\begin{bmatrix}\mathbf{x}^{1}&\cdots\mathbf{x}^{k}\end{bmatrix}\in\mathbb{C}^{N\times k}$, $\mathbf{V}=\begin{bmatrix}\mathbf{W}\mathbf{s}^{1}_{1}&\cdots\mathbf{W}\mathbf{s}^{k}_{1}\\ \vdots&\vdots&\\ \mathbf{W}\mathbf{s}^{1}_{C}&\cdots\mathbf{W}\mathbf{s}^{k}_{C}\end{bmatrix}\in\mathbb{C}^{CN\times k}$, the bi-linear measurements can be transformed into linear measurements in terms of a rank-k matrix $\textbf{U}\textbf{V}^{T}$:
@@
with a linear operator $\mathcal{A}$ that maps $\textbf{U}\textbf{V}^{T}$ to $\mathcal{F}_{\Omega}\{\Sigma_{i=1}^{k}\mathbf{s}_{j}^{i}\odot\mathbf{x}^{i}\}$. As shown in [Holme_Rosenzweig_Ong_Wilke_Lustig_Uecker_2019], the above formulation can be further relaxed to a nuclear norm minimization problem in terms of a matrix $\mathbf{Z}$,

### v3

@@ Abstract
Keywords: Self-calibration, Compressive sensing, Convex optimization, Random matrices, Parallel MRI (pMRI), Inverse problems.
@@ 1 Introduction
The inverse problem for auto-calibration is generally difficult to solve. In practice, most self-calibration algorithms rely on joint estimation of both the signals of interest and the calibration parameters using standard techniques such as maximum likelihood estimation, alternating minimization, or convex/nonconvex optimization. In this work, we extend the line of work that combines convex optimization and compression sensing in solving various auto-calibration tasks [1, 2, 3, 4]. A common scheme in such approaches includes first transforming the biconvex constraints to convex constraints using the concept of ‘lifting’ [2]; then, drawing on ideas from compressive sensing to exploit redundancy/sparsity priors in the signal model and/or the calibration model while being consistent with the convex constraints.
@@
Of particular interest is the case when $\mathbf{x}$ is transform-sparse, a characteristic often observed in real-world signals  [^1] This notion can be easily generalized to approximately transform-sparse.. That is, there is some known sparsifying transformation $\boldsymbol{\Psi}\in\mathbb{C}^{N\times N}$ such that $\boldsymbol{\Psi}\textbf{x}$ is n-sparse with $n<N$. Common choices for $\boldsymbol{\Psi}$ include the well-known DCT basis and the wavelet basis. Despite the sparsity assumption, the available measurements are still fewer than the number of unknowns (i.e., $CL\leq CN+n$). We impose an additional subspace assumption, $\mathbf{s}_{i}=\mathbf{B}\mathbf{h}_{i}$, which can be interpreted as that the calibration parameters all lie in some known subspace spanned by a tall matrix $\mathbf{B}$. Practical choices for such $\mathbf{B}$ can be a low-order polynomial basis or a low-frequency sinusoidal basis. Such choices promote smoothness in the sensing parameters [5, 6, 7, 8]. Altogether, the forward model becomes:
@@
An important observation is that Parallel Imaging (PI) studies can sometimes appear “noisy”. This is largely because the primary objective of PI techniques is to reduce imaging time, which leads to acquiring fewer data points per coil due to the need to reduce imaging time. Consequently, this reduction results in a lower signal-to-noise ratio (SNR). As a result, the SNR of PI sequences is typically lower than that of comparable non-PI sequences [9, 10].
@@
where $R$ denotes the reduction or acceleration factor for undersampling, and $g$ represents a spatially dependent term known as the geometry factor. The term $g$ relates to the number, size, and orientation of the surface coil elements and can be considered a measure of coil separation [9, 10].
@@ 2 Related work
Various auto-calibration problems of similar setups have been explored, including blind deconvolution in [3], direction-of-arrival estimation as seen in [1], blind deconvolution and demixing [4], phase retrieval via matrix completion [2], among others. The analysis of the auto-calibration problems generally differs in the following two aspects:
@@
Two primary approaches to solving the bilinear optimization problem are as follows: one involves convexifying it through methods such as linearization or lifting [3, 1, 11, 4, 2]. The other involves optimization in the natural parameter space using alternating minimization and gradient descent algorithms [12, 13, 14].
For a general discussion on the issue of injectivity and the principle of identifiability for bi-linear compressive sensing, one may refer to [15, 16].
@@ 2.1 Related work in auto-calibration and compressive sensing
However, the above nonlinear least squares problem is a challenging optimization problem since it is highly nonconvex and most of the available algorithms, such as alternating minimization and gradient descent, may suffer from easily getting trapped in some local minima.
@@
Hence, by defining the linear operator $\mathcal{A}:\mathbb{C}^{N_{1}\times N_{2}}\rightarrow\mathbb{C}^{m}$, the system of equations becomes
$y_{1},\cdots,y_{m}]=\mathcal{A}(\mathbf{X}):=\{\mathbf{b}_{i}^{*}\mathbf{X}\mathbf{a}_{i}\}_{i=1}^{m}.$ (6)
@@
An important auto-calibration problem pertaining to sparse signals arises from the task of image reconstruction using randomly coded masks [17, 18]. The forward model for this auto-calibration problem is given by:
@@
The SparseLift framework [1] solves the auto-calibration problem in (7) by combining ‘lifting’ and bi-convex compressive sensing. The authors in [1] first transform the biconvex constraint (7) to a linear constraint with respect to $\textbf{X}=\textbf{h}\textbf{x}^{T}$. Exploiting the sparsity in $\mathbf{x}$ and consequently within X, the recovery solves a convex relaxation by minimizing the $\|\mathbf{X}\|_{1}$ subject to the linear constraints. The authors show the recovery of $\mathbf{X}_{0}=\mathbf{h}_{0}\mathbf{x}_{0}^{T}$ is guaranteed with high probability when the sensing matrix $\mathbf{A}$ is either a Gaussian random matrix or with its rows chosen uniformly at random with replacement from the discrete Fourier transformation (DFT) matrix, when the number of measurements $L$ is of the order of $\mathcal{O}(kn\log^{2}(L))$. Note that the case where the signal is only approximately sparse is not analyzed. In a related work, the author in [4] extends the results of SparseLift to the case where one observes a summation of such measurements from multiple sparse signals, $\mathbf{y}=\sum\limits_{i\in[r]}\text{diag}(\mathbf{B}\mathbf{h}_{i})\mathbf{A}\mathbf{x}_{i}+\boldsymbol{\omega}$. This extension improves recovery guarantees by solving a $\ell_{1,2}$ minimization by promoting block sparsity in the lifted signals.
The idea of ‘lifting’ also applies to another class of self-calibration problems, notably the blind deconvolution problem [3, 19]. In the general form of blind deconvolution, one measures the convolution of a signal with an unknown filter. That is, $\textbf{y}=\mathbf{I}_{\Omega}(\textbf{s}*\textbf{x})+\textbf{n}$ where both s and x are unknown. Notice that the forward model for our problem in (2) can be written in the form of a convolution $\mathbf{y}=\mathbf{I}_{\Omega}(\mathbf{F}\mathbf{Bh}*\mathbf{F}\boldsymbol{\Psi^{*}}\mathbf{z})+\textbf{n}$ using the convolution theorem. When the Fourier space is fully sampled (i.e. $\mathbf{I}_{\Omega}=\mathbf{I}$ ), the problem is equivalent to recovery from
@@
where $\mathbf{w}$ is also Gaussian. The work by Ahmed et al. [3] provides a theoretically guaranteed solution to the fully sampled case in (8). Using lifting, the authors solve a nuclear norm minimization problem by exploiting low-rankness in the lifted signal. If the active coefficients of z are known and $\|\textbf{z}\|_{0}=n$, the main theorem of [3] guarantees a stable recovery when the number of measurements is, up to log factor, of the order $\mathcal{O}(n+k)$ when $\boldsymbol{\Psi}$ is a random Gaussian matrix and B is incoherent. For the more difficult case when the convolution is only subsampled, the authors in [20] provide rigorous analysis when both s and x are transform sparse (i.e., there exists $\boldsymbol{\Phi}$ and $\boldsymbol{\Theta}$ such that $\boldsymbol{\Phi}\textbf{x}$ and $\boldsymbol{\Theta}\textbf{s}$ are sparse). It provides an iterative algorithm for robust recovery guarantees with near optimal sample complexity. However, the recovery can only be guaranteed when both the sparsifying transformations (i.e. $\boldsymbol{\Phi}$ and $\boldsymbol{\Theta}$) are random Gaussians. As seen in many theories, subgaussian random matrices provide optimal measurement matrices for compressive sensing [21]. However, note that for our application case, $\boldsymbol{\Phi}=\boldsymbol{\Psi}\textbf{F}^{-1}$ and $\boldsymbol{\Theta}=\textbf{B}^{*}\textbf{F}^{-1}$, both impose structures due to physical constraints.
Similar to the parallel imaging (i.e., $C>1$ in (2)), from a practical viewpoint, a relevant scenario focuses on self-calibration from multiple snapshots. For example, blind image restoration from multiple filters [22] and self-calibration model for sensors [23, 11, 24, 25]. The goal here is to recover the unknown gains/phases $\mathbf{S}=\text{diag}(\mathbf{s})$ and a signal matrix $\mathbf{X}=[\mathbf{x}_{1},\cdots,\mathbf{x}_{p}]$ from the measurement matrix $\mathbf{Y}=[\mathbf{y}_{1},\cdots,\mathbf{y}_{p}]$ and
@@
which is a linear equation in the concatenated vector $\begin{bmatrix}\mathbf{d}\\ \mathbf{x}_{l}\end{bmatrix}$. Compared to ‘lifting’, an advantage of ‘linearization’ is that the underlying problem dimension remains unchanged. The same idea of linearization can also be found in [24, 25, 23, 26]. In [11], the authors use ‘linearization’ and solve a least-square minimization problem for (9). Furthermore, theoretical recovery guarantee is provided when $\mathbf{A}$ is Gaussian.
As observant readers already may have noticed, the ‘true dimension’ of the problem of recovering $(\textbf{h}_{1},\cdots,\textbf{h}_{c},\textbf{z})$ is $(kC+n)$. To the best of the authors’ knowledge, currently there are no convex optimization schemes which succeed with such few measurements. In this context, the work by Oymak et al. [27] should be mentioned. The authors demonstrated nonconvex procedures (e.g., minimizing a linear combination of $\ell_{0}$, $\ell_{0,2}$ or rank of the lifted matrix) that succeed with high probability, when the number of measurements is (up to log-factors) slightly larger than the information-theoretical limits. However, the theories apply only to measurement ensembles with specific properties, such as matrices with sub-Gaussian rows, subsampled standard basis (e.g. in matrix completion), or quadratic measurements for phase retrieval. Also, these procedures are not efficiently computable. For other structured measurement ensembles, it is worth noting the works by Qu et al. and Ling et al. [14, 13]. These studies address certain blind deconvolution problems with efficient algorithms and theoretical guarantees when the number of measurements is (up to log-factors) slightly larger than the information-theoretical limits. However, to the best of our knowledge, none of these methods directly apply to our measurement ensemble or require additional assumptions that do not align with our application case.
@@ 2.2.1 Relevant problem formulations and the optimization problems
According to linear algebra and the generalized sampling theorem, the maximum reduction factor is equal to the number of channels if E is of full rank. Because these conditions are rarely guaranteed in practice, the possible reduction factor is much lower. In practice, the reduction factor is typically chosen between 2 and 6 [28, 29] for parallel imaging when the number of coils is large. Thus, at the boundary case, one would expect a threshold for the maximum reduction factor and the minimum number of channels necessary to ensure the invertibility of such a linear problem.
@@
As a convex relaxation to $\ell_{0}$ minimization, the following $\ell_{1}$ minimization is often utilized for practical compressive sensing MRI applications. Since Lustig et al. [30] first demonstrated the compressive sensing approach for MRI reconstruction, it has become one of the most important tools used for modern MR imaging research. The original $\ell_{1}$ norm minimization problem solves the following convex program,
@@
where the signal x is assumed to be transform sparse. In the MRI literature, it has been demonstrated that the discrete cosine transformation, the wavelet transform, and the finite difference transform are the most commonly utilized choices for sparsifying MR images [31, 32, 33]. The choice of such $\boldsymbol{\Psi}$ is related to the field of research in sparsifying transformations [34, 35]. More recently, attention has turned to learning the sparsifying transformations from data [36, 37].
In terms of combining CS with parallel imaging, the first applications were demonstrated in [38, 39, 40, 41]. They can be seen as a direct extension of (13) to the parallel case,
$\min\limits_{\textbf{x}}\|\boldsymbol{\Psi}\textbf{x}\|_{1}+\alpha\|\textbf{x}\|_{\text{TV}}\quad\text{subject to }\textbf{E}\textbf{x}=\textbf{y},$ (14)
@@
Note that both the convex relaxations in (13) (14) and nonconvex problem in (12) assume that the coil sensitivities are known. In practice, coil sensitivities are typically determined either through a separate pre-scan [5, 42, 43] or by interpolating from a fully sampled region of k-space (Fourier space) [44, 45, 46]. However, both methods have inherent drawbacks. The pre-scan approach is susceptible to motion artifacts, environmental fluctuations, and requires additional acquisition time. The interpolation method, while faster, may introduce errors due to imperfect k-space sampling or interpolation techniques. Moreover, both approaches may necessitate the use of filtering and kernel selection to obtain reasonable estimates of the coil sensitivities, which can lead to variations in the resulting reconstructions.
@@
Prior knowledge about the object and the coil sensitivities must be incorporated to render reasonable solutions. As formulated in Joint Image Reconstruction and Sensitivity Estimation in SENSE (JSENSE) [6] and Regularized Nonlinear Inversion (NLINV) [8], the smoothness in the coil sensitivities is exploited. In JSENSE, the coil sensitivities are assumed to be expanded by a low-order polynomial basis (i.e. B) and the following bi-convex optimization problem is solved via alternating minimization,
@@
where $\mathbf{W}\in\mathbb{C}^{N\times N}$ is an invertible preconditioning matrix that penalizes high frequencies, thereby enforcing smoothness in the coil sensitivities. Other methods to enforce smoothness in the coil include expanding the coil sensitivities using low-order polynomial or sinusoidal basis functions, a method first introduced in the original SENSE paper [5] and later adopted in the analytical phantom package [7].
An extension of the NLINV (16) framework is the ENLIVE [47], which solves the following relaxed joint nonconvex optimization problem over $k$ sets of measurements,
@@
Of particular interest to us is that ENLIVE can be related to a nuclear norm minimization problem that exploits the low-rankness in a ‘lifted’ domain. Specifically, by defining the matrices
$\mathbf{y}=\begin{bmatrix}\mathbf{y}_{1}&\cdots&\mathbf{y}_{C}\end{bmatrix},\mathbf{U}=\begin{bmatrix}\mathbf{x}^{1}&\cdots\mathbf{x}^{k}\end{bmatrix}\in\mathbb{C}^{N\times k},\mathbf{V}=\begin{bmatrix}\mathbf{W}\mathbf{s}^{1}_{1}&\cdots\mathbf{W}\mathbf{s}^{k}_{1}\\ \vdots&\vdots&\\ \mathbf{W}\mathbf{s}^{1}_{C}&\cdots\mathbf{W}\mathbf{s}^{k}_{C}\end{bmatrix}\in\mathbb{C}^{CN\times k},$
the bi-linear measurements can be transformed into linear measurements in terms of a rank-k matrix $\textbf{U}\textbf{V}^{T}$:
@@
with a linear operator $\mathcal{A}$ that maps $\textbf{U}\textbf{V}^{T}$ to $\mathcal{F}_{\Omega}\{\Sigma_{i=1}^{k}\mathbf{s}_{j}^{i}\odot\mathbf{x}^{i}\}$. As shown in [47], the above formulation can be further relaxed to a nuclear norm minimization problem in terms of a matrix $\mathbf{Z}$,